# : An Open-Source SRAM Yield Analysis and Optimization Benchmark Suite

### v1

@@ Abstract
Static Random-Access Memory (SRAM) yield analysis is essential for semiconductor innovation, yet research progress faces a critical challenge: the significant disconnect between simplified academic models and complex industrial realities. The absence of open, realistic benchmarks has created a reproducibility crisis, where promising academic techniques often fail to translate to industrial practice. We present OpenYield, a comprehensive open-source ecosystem designed to address this critical gap through three core contributions: (1) A realistic SRAM circuit generator that uniquely incorporates critical second-order-effect parasitics, inter-cell leakage coupling, and peripheral circuit variations, which are typically omitted in academic studies but decisive in industrial designs. (2) A standardized evaluation platform with a simple interface and implemented baseline yield analysis algorithms, enabling fair comparisons and reproducible research. (3) A standardized SRAM optimization platform, demonstrating OpenYield’s utility in enhancing SRAM design robustness and efficiency, providing a comprehensive benchmark for optimization algorithms. OpenYield creates a foundation for meaningful academia-industry collaboration, accelerating innovation in memory design. The framework is publicly available on OpenYield:URL.
@@ I Introduction
Static Random-Access Memory (SRAM) has emerged as the dominant on-chip memory technology in modern integrated circuits, consuming up to 70% of die area in advanced AI accelerators and 60% in typical mobile System-on-Chips (SoCs) [1]. As semiconductor manufacturing pushes into sub-5nm nodes, the economic stakes of SRAM yield have become paramount—a mere 1% yield loss in high-volume production translates to millions in revenue impact, making accurate yield prediction not just a technical challenge but a critical business imperative for the $580 billion semiconductor industry [2].
@@
The computational challenge of SRAM yield analysis is daunting. Modern memory arrays require failure rates below to be viable, yet directly observing such rare events would require over Monte Carlo simulations—computationally infeasible even with today’s massive computing resources [3]. This fundamental limitation has spurred decades of research into more efficient techniques, from variance reduction methods like importance sampling [4] and statistical blockade [5], to modern machine learning approaches [6, 7] and surrogate modeling techniques [8, 9]. Each promises to dramatically reduce simulation costs while maintaining accuracy.
@@
Yet despite this rich algorithmic innovation, a troubling pattern has emerged: techniques that show exceptional promise in academic publications routinely fail when applied to industrial designs. Our comprehensive review of 45 recent SRAM yield analysis papers reveals the root cause—over 85% rely on dramatically oversimplified models that omit the very physical effects in real circuits. This systematic disconnect between academic models and industrial reality has created what we term a “reproducibility crisis” in SRAM yield analysis.
@@
The gap between academic models and industrial reality is qualitative, not merely quantitative. Real SRAM arrays exhibit complex phenomena that fundamentally alter behavior: parasitic resistances slow bitline discharge by 4-6 [10]; leakage from hundreds of unselected cells degrades read margins by 12% [11]; sense amplifier offset variations of 15-30mV shift failure boundaries unpredictably [12]; and position-dependent IR drops create systematic variations across arrays [13]. When combined, these effects create failure modes that simply cannot exist in simplified models. This modeling gap has severe practical consequences. The semiconductor industry spends $2-3 billion annually on design respins, with yield issues driving 30% of these costs. More critically, the disconnect undermines research itself: algorithms showing similar performance on simplified benchmarks may differ by orders of magnitude on real circuits; optimization techniques that appear optimal in studies fail catastrophically in practice; and without standardized, realistic benchmarks, fair comparison of approaches becomes impossible.
@@
To address this systemic challenge, we present OpenYield: an open-source ecosystem for realistic SRAM yield analysis and optimization. Our vision is to democratize access to industrial-quality SRAM modeling, enabling the academic community to develop and validate techniques on circuits that actually represent the challenges faced in modern semiconductor design. OpenYield makes three core novelties:
@@
A Python-based, scalable SRAM generator that configures any-sized array with high-quality modeling of parasitics of networks, inter-cell leakage, and peripheral circuit variations—critical effects absent in academic studies. Validation shows models without parasitics underestimate read delay and power by 18.7 and 16.7; models without peripherals underrate by 4.8 and 5.4.
@@
A Standardized Yield Analysis Platform with reference implementations of baseline algorithms (Monte Carlo, Importance Sampling variants, ML approaches) operating on realistic circuits, enabling reproducible benchmarking across different SRAM configurations.
@@
A Standardized SRAM Optimization Platform with five state-of-the-art optimization algorithms for transistor sizing under realistic conditions, enabling fair comparison and demonstrating up to 65% SNM improvements and 15% area reduction over baseline designs.
Like ImageNet [14] for computer vision or SPEC [15] for computer architecture, OpenYield establishes standardized benchmarks that enable reproducible research and accelerate innovation in memory design. The complete framework is publicly available on OpenYield:URL.
@@ II-A Physical Origins of the SRAM Yield Challenge
Process variations in semiconductor manufacturing create the fundamental yield challenge in SRAM design. As transistors shrink to atomic scales, random dopant fluctuation becomes inevitable—a 20nm transistor contains only 100 dopant atoms in its channel, making each atom’s position statistically significant [16]. These atomic-level variations manifest as threshold voltage () mismatch between transistors, with standard deviations following Pelgrom’s law: , where ranges from 2-4 mV·µm in modern processes [17].
@@ II-B Formulation of Yield Analysis and SRAM Optimization
where represents transistors’ widths (W) and lengths (L) in a SRAM cell, are objectives (e.g., power, area, noise margin), and can be yield-related constraints. Critically, these relationships change dramatically based on which physical effects the model includes.
@@ II-C The Modeling Gap: From Simplification to Crisis
Distributed parasitics create position-dependent behavior. A 256-row bitline presents fF capacitance and resistance, not as lumped elements but distributed along the wire. Cells at the array’s far end see 4-6 slower discharge rates. Simple RC models miss the incomplete swing and non-monotonic delay that distributed effects create.
Inter-cell coupling introduces data-pattern dependencies. With 255 unselected cells per bitline, cumulative leakage reaches microamps—comparable to the selected cell’s read current at low voltages [11]. The resulting margin degradation depends on stored data patterns, creating failures invisible to single-cell analysis.
Peripheral circuits often dominate yield loss. Sense amplifier (SA) offset voltage ( 20mV) directly subtracts from read margin [12]. Wordline driver skew creates systematic timing variations across columns. These effects require modeling the complete SRAM array, not just the core cell.
Yield analysis algorithms validated on simplified circuits may produce order-of-magnitude errors on real arrays. Without intervention, this gap threatens the entire research enterprise. Algorithms cannot advance without realistic validation. Industry cannot adopt techniques proven only on toy models. The field needs standardized, realistic benchmarks—exactly what OpenYield provides.
@@ II-D SOTA SRAM Tools and Their Limitations
OpenRAM [18] compiles complete SRAMs from specification to GDS with diverse configurations and manufacturable layouts, but lacks variation modeling and yield analysis infrastructure.
SRAM22 [19] provides automated embedded SRAM generation with compact layouts, but similarly offers no statistical analysis or yield optimization features.
AnalogGym [20] benchmarks general analog synthesis algorithms but cannot capture SRAM-specific phenomena like massive parallelism, statistical extremes, and critical layout dependencies.
Commercial tools accurately model these effects but remain inaccessible due to proprietary licenses, black-box implementations, and high costs that exclude academic research. OpenYield addresses this gap by combining industrial-quality models with open-source accessibility, enabling reproducible research and advancing SRAM yield analysis.
@@
The OpenYield ecosystem follows a modular architecture depicted in Fig. missing 1. The core SRAM Circuit Generator produces PySpice-based netlists [21] that incorporate user configurations and realistic physical effects. These netlists are analyzed by an engine using open-source SPICE tools like Xyce [22].
Two principal blocks leverage the simulation insights: Yield Analysis Algorithms (Sect. missing III-B) implement statistical techniques for yield estimation and failure probability assessment, while the Design Optimization Framework (Sect. missing III-C) automatically adjusts design parameters (e.g., bitcell transistor sizes) to enhance SRAM robustness and performance.
Users interact with OpenYield by providing comprehensive inputs including SRAM configurations (arbitrary array dimensions, memory cell types from 6T standard with 8T/10T provisions), technology characteristics (currently FreePDK45 [23] support), design parameters, process variation statistics (local intra-die random variations applied independently to each transistor using Gaussian distributions), and optimization goals for Monte Carlo simulations. The platform delivers multiple outputs: SPICE netlists (.sp format) ensuring broad simulator compatibility, detailed yield reports, failure probability estimates, characterized performance metrics, and optimized design parameters. Monte Carlo parameterization can be embedded directly in model cards or exported as data tables for importance sampling-based yield analysis algorithms. The platform follows open-source principles, encouraging community contributions and extensions.
@@ III-A Hierarchical SRAM Generator
The cornerstone of OpenYield is its novel SRAM circuit generator, engineered to produce unified benchmark circuits that transcend common academic simplifications and approach industrial SRAM complexities.
@@ III-A1 Circuit Architecture
As illustrated by Fig. missing 2, the 6T SRAM cell comprises cross-coupled inverters (M0-M1, M4-M5) for storage and PG (M2-M3) for read/write control. Process variations, particularly random dopant fluctuation, create threshold voltage mismatches that drive yield loss through specific failure mechanisms:
Read Upsets occur when bitline BLB discharge through M3-M1 causes node V(QB) of M4-M0 inverter, destructively flipping the cell state. Read Static Noise Margin (RSNM) quantifies the cell’s immunity to this—insufficient RSNM (typically 50mV) directly causes the failure.
@@
All these failure mechanisms can be simulated and analyzed through OpenYield. The generator constructs complete array hierarchies including: core cells in configurable row/column arrangements, wordline drivers sized for distributed loads, column multiplexers with shared sense amplifiers, precharge/write driver circuits, sense amplifiers with realistic offset distributions, and configurable control logic. This comprehensive approach distinguishes OpenYield from existing academic generators that produce idealized circuits.
@@ III-A2 Second-Order Physical Effects Modeling
OpenYield integrates critical physical effects that profoundly influence yield in advanced technologies:
Parasitic Network Generation: Distributed -shaped RC segments model vital conductive paths (bitlines, wordlines) rather than oversimplified lumped capacitances. Users can specify resistance/capacitance values for each subcircuit, with defaults based on typical SRAM layout topologies.
Inter-Cell Leakage Coupling: Models various leakage mechanisms (subthreshold, gate leakage) to simulate neighboring cell impacts on actively accessed cells—particularly significant under low VDD conditions.
Peripheral Circuit Variation: Instantiates realistic models for sense amplifiers (with systematic/random offset voltage distributions), wordline drivers (with delay variations and skews), and precharge/write drivers (with voltage level and timing variations). These peripheral variations are modeled conjunctively with core cell variations for holistic yield assessment.
@@ III-B1 Monte Carlo Reference
Standard Monte Carlo serves as the accuracy baseline, automating variation sampling, SPICE simulation, and pass/fail evaluation. Enhanced with stratified sampling and control variates, the implementation reduces required samples by - while maintaining unbiased estimation—essential for validating advanced methods.
@@ III-B2 Importance Sampling Variants
Four IS methods address different yield analysis challenges. Mean-shifted IS (MNIS) [4] translates distributions toward the most probable failure point: subject to failure boundary . This achieves speedup when single modes dominate, particularly for read failures driven by sense amplifier mismatch. Adaptive Compressed Sampling (ACS) [24] builds sparse failure representations using compressed sensing, reducing samples by 90% for smooth failure boundaries. Adaptive IS (AIS) [25] iteratively refines proposal distributions through cross-entropy minimization, automatically discovering multiple failure modes. High-dimensional Sparse Compressed Sampling (HSCS) [26] exploits sparsity to scale efficiently to hundreds of parameters, crucial for full-array analysis with systematic variations.

### v2

@@ Abstract
Static Random-Access Memory (SRAM) yield analysis is essential for semiconductor innovation, yet research progress faces a critical challenge: the large gap between simplified academic models and the complexities observed in practice. The lack of open, higher-fidelity benchmarks has hindered reproducibility and transferability, as promising academic techniques often fail to carry over to more realistic settings. We present OpenYield, an open-source ecosystem that aims to narrow this gap through three contributions: (i) An SRAM circuit generator that explicitly incorporates second-order effects (interconnect/line parasitics, inter-cell leakage coupling, and peripheral-circuit variations) that are commonly omitted in academic studies. (ii) A standardized evaluation platform with a simple interface and baseline yield-analysis implementations to enable fair comparisons and reproducible research on these higher-fidelity circuits. (iii) An optimization platform for transistor-level sizing under these models, supporting reproducible studies of robustness/efficiency trade-offs. OpenYield aims to foster more reproducible and transferable progress in SRAM-yield research. The framework is publicly available at OpenYield:URL.
@@ I Introduction
Static Random-Access Memory (SRAM) has emerged as the dominant on-chip memory technology in modern integrated circuits, consuming up to 70% of die area in advanced AI accelerators and 60% in typical mobile System-on-Chips (SoCs) [chandrakasan2019sram]. As semiconductor manufacturing pushes into sub-5nm nodes, the economic stakes of SRAM yield have become paramount—a mere 1% yield loss in high-volume production translates to millions in revenue impact, making accurate yield prediction not just a technical challenge but a critical business imperative for the $580 billion semiconductor industry [agarwal2008statistical].
@@
The computational challenge of SRAM yield analysis is daunting. Modern memory arrays require failure rates below to be viable, yet directly observing such rare events would require over Monte Carlo simulations—computationally infeasible even with today’s massive computing resources [kanj2006mixture]. This fundamental limitation has spurred decades of research into more efficient techniques, from variance reduction methods like importance sampling [dolecek2008breaking] and statistical blockade [singhee2008practical], to modern machine learning approaches [bouhlila2024machine, yoon2022machine] and surrogate modeling techniques [wang2017efficient, yao2014efficient]. Each promises to dramatically reduce simulation costs while maintaining accuracy.
@@
Yet despite this rich algorithmic innovation, a troubling pattern has emerged: techniques that show exceptional promise in academic publications often fail to transfer to production designs. Our comprehensive review of 45 recent SRAM yield analysis papers reveals the root cause—over 85% rely on dramatically oversimplified models that omit the very physical effects in real circuits. This systematic disconnect between academic models and production reality has created what we term a “reproducibility crisis” in SRAM yield analysis.
@@
The gap between academic models and industrial reality is qualitative, not merely quantitative. Real SRAM arrays exhibit complex phenomena that fundamentally alter behavior: parasitic resistances and capacitances slow bitline discharge by 4-6 [shen2024deep]; leakage from hundreds of unselected cells degrades read margins by 12% [shen2024ultra8t]; sense amplifier offset variations of 15-30mV shift failure boundaries unpredictably [abu2011characterization, shen2022single]; and position-dependent IR drops create systematic variations across arrays [shen2024sram]. When combined, these effects create failure modes that simply cannot exist in simplified models. This modeling gap has severe practical consequences. The semiconductor industry spends $2-3 billion annually on design respins, with yield issues driving 30% of these costs. More critically, the disconnect undermines research itself: algorithms showing similar performance on simplified benchmarks may differ by orders of magnitude on real circuits; optimization techniques that appear optimal in studies fail catastrophically in practice; and without standardized, higher-fidelity benchmarks, fair comparison of approaches becomes impossible.
@@
To address this systemic challenge, we present OpenYield: an open-source ecosystem for SRAM yield analysis and optimization that aims to narrow the gap between research circuits and real SRAM modules in the on-chip systems. Our goal is to make higher-fidelity, open benchmarks more accessible, enabling the academic community to develop and validate techniques on circuits that incorporate key second-order effects observed in practice. OpenYield makes three core novelties:
@@
A Python-based, scalable SRAM generator that configures any-sized array and explicitly models key wire parasitics, inter-cell leakage, and peripheral circuit variations—effects often omitted in academic studies. Validation shows models without parasitics underestimate read delay and power by 3.7 and 2.5; models without peripherals underrate by 2.3 and 1.28.
@@
A Standardized Yield Analysis Platform with reference implementations of baseline algorithms (Monte Carlo and importance sampling variants) operating on circuits that include these effects, enabling reproducible benchmarking across different SRAM configurations.
@@
A Standardized SRAM Optimization Platform with five state-of-the-art optimization algorithms for transistor sizing under these models, enabling fair comparison and demonstrating up to 65% SNM improvements and 15% area reduction over baseline designs.
OpenYield aims to transform SRAM yield analysis/optimization research by providing higher-fidelity test circuits and baseline methods in an open-source platform. Like ImageNet [Deng2009imagenet] for computer vision or SPEC for computer architecture [Henning2006spec], it establishes standardized benchmarks that enable reproducible research and new technical directions. The complete framework, with documentation and examples, will be publicly available to maximize adoption.
@@ II-A Physical Origins of the SRAM Yield Challenge
Process variations in semiconductor manufacturing create the fundamental yield challenge in SRAM design. As transistors shrink to atomic scales, random dopant fluctuation becomes inevitable—a 20nm transistor contains only 100 dopant atoms in its channel, making each atom’s position statistically significant [asenov2003random]. These atomic-level variations manifest as threshold voltage () mismatch between transistors, with standard deviations following Pelgrom’s law: , where ranges from 2-4 mV·µm in modern processes [pelgrom1989matching].
@@ II-B Formulation of Yield Analysis and SRAM Optimization
where represents transistors’ widths (W) and lengths (L) in a SRAM cell, are objectives (e.g., power, area, noise margin), and can be yield-related timing constraints. Critically, these relationships change dramatically based on which physical effects the model includes.
@@ II-C The Modeling Gap: From Simplification to Crisis
Distributed parasitics create position-dependent behavior. A 256-row bitline presents fF capacitance and resistance, not as lumped elements but distributed along the wire. Cells at the array’s far end see 4-6 slower discharge rates [shen2024deep]. Simple RC models miss the incomplete swing and non-monotonic delay that distributed effects create.
Inter-cell coupling introduces data-pattern dependencies. With 255 unselected cells per bitline, cumulative leakage reaches microamps—comparable to the selected cell’s read current at low voltages [shen2024ultra8t]. The resulting margin degradation depends on stored data patterns, creating failures invisible to single-cell analysis.
Peripheral circuits often dominate yield loss. Sense amplifier (SA) offset voltage ( 20mV) directly subtracts from read margin [abu2011characterization, shen2022single]. Wordline driver skew creates systematic timing variations across columns. These effects require modeling the complete SRAM array, not just the core cell.
Other second-order effects include: (i) layout-aware spatial/systematic variation (pattern density, well proximity, strain/stress, orientation); (ii) supply-network effects (static IR drop and dynamic supply bounce on VDD/VSS); and (iii) temperature and aging (self-heating/thermal gradients, BTI/HCI-induced shifts).
We prioritize the first three effects because, in mainstream 2D CMOS SRAMs and the low-voltage operating points we study, these are the dominant contributors to read/write margin loss and sense errors. Other effects remain important and are compatible with our framework, and we leave the integration of them to future work.
@@ II-D SOTA SRAM Tools and Their Limitations
Current SRAM tools serve specific niches but fail to address yield analysis needs comprehensively. OpenRAM [guthaus2016openram] compiles complete SRAMs from specification to GDS with diverse configurations and manufacturable layouts, but lacks variation modeling and yield analysis infrastructure. SRAM22 [SRAM22] provides automated embedded SRAM generation with compact layouts, but similarly offers no statistical analysis or yield optimization features. AnalogGym [li2024analoggym] benchmarks general analog synthesis algorithms but cannot capture SRAM-specific phenomena like massive parallelism, statistical extremes, and critical layout dependencies.
Commercial tools accurately model these effects but remain inaccessible due to proprietary licenses, black-box implementations, and high costs that exclude academic research. OpenYield addresses this gap by providing open-source models that incorporate many second-order effects beyond common academic simplifications, enabling reproducible research and more informative comparisons.
@@
The OpenYield ecosystem follows a modular architecture depicted in Fig. 1. The core SRAM Circuit Generator produces PySpice-based netlists [PySpice] that incorporate user configurations and relevant physical effects. These netlists are analyzed by an engine using open-source SPICE tools like Xyce [keiter2013xyce].
Two principal blocks leverage the simulation insights: Yield Analysis Algorithms (Sect. III-B) implement statistical techniques for yield estimation and failure probability assessment, while the Design Optimization Framework (Sect. III-C) automatically adjusts design parameters (e.g., bitcell transistor sizes) to enhance SRAM robustness and performance.
Users interact with OpenYield by providing comprehensive inputs including SRAM configurations (arbitrary array dimensions, memory cell types from 6T standard with 8T/10T provisions), technology characteristics (currently FreePDK45 [FreePDK45] support), design parameters, process variation statistics (local intra-die random variations applied independently to each transistor using Gaussian distributions and global process corners), and optimization goals for Monte Carlo simulations. The platform delivers multiple outputs: SPICE netlists (.sp format) ensuring broad simulator compatibility, detailed yield reports, failure probability estimates, characterized performance metrics, and optimized design parameters. Monte Carlo parameterization can be embedded directly in model cards or exported as data tables for importance sampling-based yield analysis algorithms. The platform follows open-source principles, encouraging community contributions and extensions.
@@ III-A Hierarchical SRAM Generator
The cornerstone of OpenYield is its novel SRAM circuit generator, engineered to produce unified benchmark circuits that transcend common academic simplifications and approach SRAM complexities observed in practice.
@@ III-A1 Circuit Architecture
As illustrated by Fig. 2, the 6T SRAM cell comprises cross-coupled inverters (M0-M1, M4-M5) for storage and PG (M2-M3) for read/write control. Process variations, particularly random dopant fluctuation, create threshold voltage mismatches that drive yield loss through specific failure mechanisms:
Read Upsets occur when bitline BLB discharge through M3-M1 causes node of M4-M0 inverter, destructively flipping the cell state. Read Static Noise Margin (RSNM) quantifies the cell’s immunity to this—insufficient RSNM (typically 50mV) directly causes the failure.
@@
All these failure mechanisms can be simulated and analyzed through OpenYield. The generator constructs complete array hierarchies including: core cells in configurable row/column arrangements, wordline drivers sized for distributed loads, column multiplexers with shared sense amplifiers, precharge/write driver circuits, sense amplifiers with statistical offset distributions, and configurable control logic. This comprehensive approach distinguishes OpenYield from existing academic generators that produce idealized circuits.
@@ III-A2 Second-Order Physical Effects Modeling
OpenYield integrates critical physical effects that profoundly influence yield in advanced technologies, as listed below.
Parasitic Network Generation: Distributed -shaped RC segments model vital conductive paths (bitlines, wordlines) rather than oversimplified lumped capacitances. Users can specify resistance/capacitance values and segment numbers for each subcircuit, with defaults based on typical SRAM layout topologies.
Inter-Cell Leakage Coupling: Models various leakage mechanisms (subthreshold, gate leakage) to simulate neighboring cell and peripherals impacts on actively accessed cells—particularly significant under low VDD conditions and deeper SRAM designs.
Peripheral Circuit Variation: Instantiates statistical models for decoders (capturing stage-dependent delay on the critical access path and output slew that feeds into the wordline drivers), wordline drivers (with delay variations and skews, and sensitivity to decoder output slew), sense amplifiers (with systematic/random offset voltage distributions), and precharge/write drivers (with voltage level and timing variations). These peripheral variations are modeled conjunctively with core cell variations for holistic yield assessment.
Other second-order effects discussed in Sect. II-C—including power-grid IR drop, layout-aware spatial/systematic variation driven by device geometry, long-term aging (BTI/HCI), and thermal gradients—are not modeled by default. The framework provides hooks to incorporate them: IR drop can be incorporated by instantiating additional RC elements on the supply rails to model impedance; layout-aware variation can be parameterized via physical design attributes (e.g., diffusion areas/perimeters) to inform device-mismatch models; and aging/thermal effects are typically secondary at our scales and can be modeled as additional parametric variations.
@@ III-B1 Monte Carlo Reference
Standard Monte Carlo is widely regarded as the “gold standard” for yield estimation due to its unbiased nature and guaranteed convergence as the number of samples increases, albeit at the cost of high computational effort for rare-event problems [shi2018fast, xing2024every, dolecek2008breaking]. This method serves as the accuracy baseline, automating variation sampling, SPICE simulation, and pass/fail evaluation. Enhanced with stratified sampling and control variates, the implementation reduces required samples by - while maintaining unbiased estimation—essential for validating advanced methods.
@@ III-B2 Importance Sampling Variants
Importance Sampling (IS) and its variants are considered robust alternatives for yield analysis: they require fewer parameters and hyperparameters than surrogate-based approaches, and they reliably converge for a wide range of yield estimation problems [dolecek2008breaking, shi2018fast, xing2024every]. Four IS methods address different yield analysis challenges. Mean-shifted IS (MNIS) [dolecek2008breaking] translates distributions toward the most probable failure point: subject to failure boundary . This achieves speedup when single modes dominate, particularly for read failures driven by sense amplifier mismatch. Adaptive Compressed Sampling (ACS) [shi2019adaptive] builds sparse failure representations using compressed sensing, reducing samples by 90% for smooth failure boundaries. Adaptive IS (AIS) [shi2018fast] iteratively refines proposal distributions through cross-entropy minimization, automatically discovering multiple failure modes. High-dimensional Sparse Compressed Sampling (HSCS) [wu2016hyperspherical] exploits sparsity to scale efficiently to hundreds of parameters, crucial for full-array analysis with systematic variations.
Surrogate model-based methods [yin2022efficient] are not considered here, as they typically demand significant model training and hyperparameter tuning, making them more suitable for extremely complex, high-dimensional problems where MC is infeasible and IS-based methods may also fail.